# A Study on Asynchronous Vote-based Blockchains

### v1

@@ 1 Introduction
Traditionally, vote-based blockchains use binary agreement, where nodes either vote to accept or reject a block, typically proposed by a periodically elected leader. They require the existence of at most one block for each voting instance. The honest nodes must make the same decision for the block in order to maintain consistency. They are commonly used in blockchain implementations, such as Tendermint [6], Algorand [15], and HotStuff [26]. However, in the asynchronous communication model, where no time-bound assumptions are made, it becomes impossible to discern whether a leader is an adversary (i.e., chooses not to share its block) or just being slow (i.e., other nodes do not receive its block in time). Setting a time-out to replace a silent leader is not a solution as it contradicts the communication model. Thus, these protocols cannot guarantee liveness in consensus derivation in asynchronous settings as they may not terminate.
Asynchronous Common Subset (ACS) based blockchain protocols [11, 2, 21, 14] address the liveness issue by assembling a block for advancing a blockchain using transaction sets submitted by at least $N-f$ nodes in an $N$-node system with $f$ potential adversaries. Liveness is ensured as honest messages will eventually be delivered. ACS protocols usually require $O(N^{3})$ messages due to using Multi-valued Validated Byzantine Agreement (MVBA) protocols or concurrently using multiple instances of Asynchronous Binary Agreement (ABA) protocols. Moreover, they can lead to transaction overlaps and greatly increase the protocol’s cost, even with a small number of distinct transactions.
@@
Moreover, the honest nodes do not need to vote for the same block per voting round. As such, an honest node does not need to wait until it receives $f+1$ acknowledgments of a block before voting for this block. This design allows leaders to broadcast blocks without the need for receivers to broadcast acknowledgments, making our design the first one to use only a message and only $O(N)$ message complexity for a view change, even without threshold signatures [^1] HotStuff [26] uses threshold signatures to achieve $O(N)$ message complexity for view changes. Message complexity [8] is the (expected) number of messages sent to the entire network. E.g., if a node sends a message of its own to all the $N$ nodes, the message complexity would be $O(N)$..
@@
Instant-Runoff Voting: It eliminates less supported blocks round-by-round until one block remains. This not only gradually aligns honest nodes but also gradually limits voting options for adversaries. When only one block remains, a silent vote implies acquiescence rather than disagreement.
@@
Reduced Complexity: Our protocol significantly decreases the message complexity. It achieves an $O(m+\theta N)$ message count and requires a message complexity of at most $O(mN+\theta N^{2})$ if each node sends its vote directly to every other node, where $m$ represents the number of blocks and $\theta$ is the number of voting rounds used. If implemented in a manner similar to HotStuff, where the leader collects votes, the total message is reduced to at most $O(mN+\theta N)$. In a standard asynchronous network with arbitrary but finite network delays, our protocol uses at least two rounds of voting before the consensus is established. This is consistent with binary agreement protocol HotStuff-2 [20], which operates only in synchronous or partially synchronous environments. Our protocol requires an amortized $\log_{2}(N)+\log_{2}(m)$ voting rounds, only in unrealistic scenarios where the adversary deterministically controls the message delays of all nodes.
@@
2 Related Works
Here, we review the existing asynchronous blockchain protocols and the strong BFT consensus protocols.
2.1 ACS-based Protocols
2.2 DAG-based Protocols
Direct Acyclic Graph (DAG)-based BFT protocols [18, 10, 25] use DAGs to transmit transactions. Each node sends a block linked to $N-f$ blocks from the preceding height. A leader block is selected periodically to order and finalize the linked blocks. But, to avoid equivocation, they must use reliable broadcast for each block, driving the overall message complexity to $O(N^{3})$.
2.3 Strong BFT Consensus Protocols
Safety. An adversary may double vote or send its votes only to a subset of honest nodes. However, this cannot trick some nodes into believing that an input value has been advanced to a new round when it actually did not. This is because less than $N-f$ nodes will confirm that the input value has received more than $f$ votes. There could be different sets of $N-f$ reports, resulting in some nodes considering that an input value was not confirmed, even though it was confirmed. But, these nodes will work this out eventually during the third phase as there must be more than $f$ nodes who have accepted this input value into the next round. So, after each voting round, the honest nodes reach the same set of remaining input values.
@@ 3.1 Assumptions on System Model
General Assumptions. We assume the availability of hash functions, digital signatures, and a public-key infrastructure (PKI). The system operates with a predetermined $N$ number of nodes. A node identity table is shared among all participating nodes.
Strong Adaptive Adversary. A Byzantine adversary comprises $f\leq\lfloor(N-1)/3\rfloor$ nodes, capable of deviating from the standard protocol through various means. Once an adversary attempts to corrupt a node, the node is immediately corrupted. Moreover, messages sent from the node before corruption but not yet delivered can be erased (or clawed back) by the adversary.
@@
Network Assumptions. Our framework adopts the standard asynchronous communication assumption, where a message sent from one honest node to another eventually reaches its destination, albeit subject to arbitrary yet finite communication delays. We support any network topology as long as this assumption is held.
@@ 3.2 Validated Strong BFT Consensus
Conversely, a blockchain validates a block $B$ by reconciling $B$ and the previous blocks of $B$, which forms $\pi_{B}$. Honest nodes abstain from committing $B$ if $Q(B,\pi_{B})=\text{False}$, regardless of how many nodes have voted for it. As honest nodes will only vote for a correct block, and if a protocol can assist all honest nodes to eventually converge to the same $B^{\prime}$ before committing, where $Q(B^{\prime},\pi_{B^{\prime}})=\text{True}$, then validated strong BFT consensus can tolerate $f\leq\lfloor(N-1)/2\rfloor$ and $f\leq\lfloor(N-1)/3\rfloor$ for synchronous and asynchronous settings respectively, a clear security advantage over strong BFT consensus.
@@
Challenges in guaranteeing liveness of consensus derivation with the leader-based design The strong BFT consensus protocol in Sec. 2.3 requires a known set of $m$ input values and tolerates only a small $f$. It does not consider the origin of the input values. If $m$ leaders propose them, then consensus liveness in asynchronous settings is not guaranteed unless $m\geq N-f$. Also, when $m>2$ and $f\leq\lfloor(N-1)/3\rfloor$, it is possible that no input value receives more than $f$ votes, making it impossible to discard a block that doesn’t receive $f+1$ votes.
To address these issues, we propose a simple protocol that does not require a known $m$ but ensures that, via multiple voting rounds, honest nodes will converge to the same value. In the first voting round, nodes vote for the first valid input value they received. In subsequent rounds, they vote for the most supported one based on their knowledge. But, this design has potential liveness issues: (1) adversaries may send conflicting votes to different nodes, causing them to fail to converge to the same value; (2) random message delays can cause nodes that converged on a value in one round to disperse to different values in the next round.
In the next section, we describe this protocol in more detail and how we ensure liveness in establishing consensus.
@@ 4.1 Key Elements of the Protocol
To help understand our proposal, we first highlight five key aspects of our protocol that depart from existing asynchronous protocols for achieving strong BFT consensus:
1). Periodic Leader-based Block Proposal and Allowing Unknown $m$: The classical approach doesn’t specify the origin of the $m$ input values. Implementing it in an asynchronous leader-based blockchain would require $m\geq N-f$ nodes to propose blocks. In contrast, our protocol uses a single periodically rotating leader node to propose a block at each time window. A block can arrive at any time due to arbitrary but finite message delays. We guarantee consensus even with new blocks being added to the voting pool during voting rounds.
2). Increased Adversarial Tolerance: The classical approach tolerates $f\leq\lfloor(N-1)/(m+1)\rfloor$ adversarial nodes, $m>2$. Our protocol allows $f\leq\lfloor(N-1)/3\rfloor$.
3). Forcing Verdict Unification via Voting Rules: The structure of votes is designed to support verification of voting rule compliance. It enables nodes to gradually align on a single block and forces adversary nodes to comply with the voting rules as well.
4). Concurrent Consensus Derivation for Multiple Instances: In contrast to the classical approach, which runs only one instance at a time, our protocol nests and concurrently executes multiple contingent instances.
5). Concurrent Voting Rounds: The classical approach requires receiving $N-f$ votes before one starts the next voting round. We introduce flexibility by the relaxed design detailed in Sec. 4.7.2. Nodes can begin the next round without waiting for votes from others.
@@ 4.2 Definitions
Vote Chain. The vote chain for node $i$ consists of successive votes, where each vote at height $k\geq 1$ has a directed link to node $i$’s vote at height $k-1$. All votes at height 00 are cast on the Genesis block.
@@
Vote ($V^{(k)}_{i}$). This is the vote by node $i$ for a block of consensus height $k$. For $k\geq 1$, each vote of vote height $k$ must link to $N-f$ votes at vote height $k-1$. These links are created by attaching their hash values to $V^{(k)}_{i}$.
@@
Vote View ($VV_{i}^{(k)}$). This is an array of integers, where $VV_{i}^{(k)}[B]$ contains the sum of the votes cast on $B$ in $G_{i}^{(k)}$. Note that $VV_{i}^{(k)}$ does not count $V^{(k)}_{i}$.
Vote Count ($VC_{i}^{(k)}$). This is an array of integers, where $VC_{i}^{(k)}[B]$ sums the votes cast on block $B$ and any blocks up until consensus height $k-1$ stemming from $B$, using the data from $VV_{i}^{(k)}$.
@@ 4.3 Consensus Chain Propagation
For simplicity of explanation, we assume all $N$ nodes are indexed and have access to a shared clock [^2] For example, there can be a pre-set agreement for nodes to start the system together at 00:00 UTC on a day., denoted by $T$, which is not required to align perfectly among the nodes. Node 0 generates the Genesis block at $T=0$, which is consensually accepted by all nodes. At the beginning of each predefined period $\Delta BI$, the node with index $i=\lfloor\frac{T}{\Delta BI}\rfloor\mod N$ proposes a block to extend the consensus chain known to node $i$  [^3] Branches (forkings) may exist, resulting in multiple consensus chains. Because of arbitrary but finite network delays, these consensus chains may or may not be known to node $i$.. As nodes are not assumed to be aligned precisely according to $T$, an adversary could effectively propose a block at any time for any branch. Fig. 1 shows an example of consensus chain propagation. All blocks stem from the Genesis block and form a tree structure.
@@ 4.4 Voting
All votes at height 0 are for the Genesis block. An honest node $i$ casts a vote at vote height $k>0$ only upon receipt of $N-f$ votes of height $k-1$  [^4] Note that if some of the $N-f$ votes are labeled as non-compliant votes by the node, this node will wait until it receives $N-f$ compliant votes. Sec. 4.4.1 discusses when a vote will be labeled non-compliant., which can be used to construct $G_{i}^{(k)}$ and derive both $VV_{i}^{(k)}$ and $VC_{i}^{(k)}$  [^5] Note that $G_{i}^{(k)}$ may contain double-voting votes, which need to be deducted before deriving $VC_{i}^{(k)}$. See the filtering rule in Sec. 4.4.1..
@@
      Let $B^{\prime}_{k^{\prime}}$ be the block at consensus height $k^{\prime}$ that $B_{k}$ stems from, where $0\leq k^{\prime}<k$, and $B^{\prime}_{0}$ be the Genesis block. $\forall k^{\prime}\in\{1,...,k-1\}$, $B^{\prime}_{k^{\prime}}$ must have the highest vote count among blocks at consensus height $k^{\prime}$ that stem from $B^{\prime}_{k^{\prime}-1}$ according to $VC_{i}^{(k)}$.
@@
Figure 1: An example of consensus chain propagation: Node 1 generated Block 1, stemming from Block 0. Block 2 did not extend Block 1; instead, it extended Block 0. Block 3 extended Block 1, etc. This figure also illustrates the vote view $VV_{j}^{(3)}$ and the vote count $VC_{j}^{(3)}$. The numbers in the blue and red circles on Block X are $VV_{j}^{(3)}[X]$ and $VC_{j}^{(3)}[X]$ respectively. $G_{j}^{(3)}$ contains votes until height 2 to the knowledge of node $j$.
@@
4:         Generate $V_{i}^{(currentHeight)}$ that links to $N-f$ compliant votes of $currentHeight-1$.
5:         Cast $V_{i}^{(currentHeight)}$ on a block of $currentHeight$ from the most supported branch (per criteria. 2 of the honest model) after filtering out double-voting votes.
@@
Vote reasoning. When receiving $V_{j}^{(k)}$ from any node $j$, any node can acquire the blocks known to node $j$ and their associated votes via $G_{j}^{(k)}$. Fig. 1 shows an example with six blocks up to height 3 and the information derived from $V_{j}^{(3)}$. The numbers in the blue and red circles on block $X$ are $VV_{j}^{(3)}[X]$ and $VC_{j}^{(3)}[X]$ respectively. No information about block 4 and block 6 are indicated in $V_{j}^{(3)}$ as $G_{j}^{(3)}$ only contains votes up until height 2. One cannot infer if node $j$ knows any block of height 3 base on $V_{j}^{(3)}$. But, $V_{j}^{(3)}$ must be cast for block 6 as it is in the branch rooted by block 5, which has the highest vote count among blocks of height 2 stemming from block 1. Block 1 has the highest vote count among blocks of height 1 that stems from the Genesis block.
Remarks: The reason for using $VC$, which sums up the votes of a block’s branch as the vote count of this block, is that once this block is consensually accepted, later consensually accepted blocks must come from its branch. Thus, the goal is to decide on which branch (rooted by a block) is most supported, not merely the most supported block.

### v2

# Validated Strong Consensus Protocol for Asynchronous Vote-based Blockchains

@@ 1 Introduction
Traditionally, vote-based blockchains use binary agreement, where nodes either vote to accept or reject a block, typically proposed by a periodically elected leader. They require the existence of at most one block for each voting instance. The honest nodes must make the same decision for the block in order to maintain consistency. They are commonly used in blockchain implementations, such as Tendermint [5], Algorand [16], and HotStuff [27]. However, in the asynchronous communication model, where no time-bound assumptions are made, it becomes impossible to discern whether a leader is controlled by the adversary (i.e., chooses not to share its block) or just being slow (i.e., other nodes do not receive its block in time). Setting a time-out to replace a silent leader is not a solution as it contradicts the communication model. Thus, these protocols cannot guarantee liveness in consensus derivation in asynchronous settings as they may not terminate.
Asynchronous Common Subset (ACS) based blockchain protocols [2, 22, 15] address the liveness issue by assembling a block for advancing a blockchain using transaction sets submitted by at least $N-f$ nodes in an $N$-node system with $f$ potential Byzantine nodes. Liveness is ensured as honest messages will eventually be delivered. ACS-based protocols usually require $O(N^{3})$ messages due to using Multi-valued Validated Byzantine Agreement (MVBA) protocols or concurrently using multiple instances of Asynchronous Binary Agreement (ABA) protocols. Moreover, they can lead to transaction overlaps and greatly increase the protocol’s cost, even with a small number of distinct transactions.
@@
Moreover, honest nodes do not need to vote for the same block in each voting round. As such, an honest node does not need to wait until it receives $N-f$ acknowledgments for the same view number before voting for a block in that view. This design allows a leader node to broadcast blocks without requiring $N-f$ nodes to recognize it as the leader first, making our design the first to use only a single message and $O(N)$ message complexity for a view change, even without threshold signatures [^1] HotStuff [27] uses threshold signatures to achieve $O(N)$ message complexity for view changes. Message complexity [8] is the (expected) number of messages sent to the network. E.g., if a node sends a message of its own to all the $N$ nodes, the message complexity would be $O(N)$..
@@
Instant-Runoff Voting: It eliminates less supported blocks round-by-round until one block remains. This not only gradually aligns honest nodes but also gradually limits voting options for Byzantine nodes. When only one block remains, a silent vote implies acquiescence rather than disagreement.
@@
Reduced Complexity: Our protocol significantly decreases the message complexity. It achieves an $O(m+\theta N)$ message count and requires a message complexity of $O(mN+\theta N^{2})$ if each node sends its vote directly to every other node, where $m$ represents the number of blocks and $\theta$ is the number of voting rounds used. If implemented in a manner similar to HotStuff, where the leader collects votes, the total message is reduced to at most $O(mN+\theta N)$. In a standard asynchronous network with arbitrary but finite network delays, our protocol can use as small as two rounds of voting in the best case before the consensus is established . This is consistent with binary agreement protocol HotStuff-2 [21], which operates only in synchronous or partially synchronous environments. Our protocol requires an amortized $\log_{2}(mN)$ voting rounds, only in unrealistic scenarios where the adversary deterministically controls the message delays of all nodes.
@@ 2.1 Assumptions on System Model
General Assumptions. We assume the availability of hash functions, digital signatures, and a public-key infrastructure (PKI). The system operates with a preset $N$ number of nodes.
Strong Adaptive Adversary. A adversary comprises $f\leq\lfloor(N-1)/3\rfloor$ Byzantine nodes, capable of deviating from the standard protocol through various means. Once the adversary attempts to corrupt a node, the node is immediately corrupted. Moreover, messages sent from the node before corruption but not yet delivered can be erased (or clawed back) by the adversary.
@@
Network Assumptions. Our framework adopts the standard asynchronous communication assumption, where a message sent from one honest node to another eventually reaches its destination, albeit subject to arbitrary, unknown, yet finite communication delays. We support any network topology as long as this assumption is held.
@@ 2.2 Validated Strong BFT Consensus
In contrast, a blockchain validates a block $B$ (if $B$ contains transactions that contradict previous blocks) by reconciling $B$ and the previous blocks of $B$, which form $\pi_{B}$. Honest nodes abstain from committing $B$ if $Q(B,\pi_{B})$ is false, regardless of how many nodes have voted for it. As honest nodes will only vote for a correct block, and if a protocol can assist all honest nodes to eventually converge to the same $B^{\prime}$ before committing, where $Q(B^{\prime},\pi_{B^{\prime}})$ holds true, then validated strong BFT consensus can tolerate $f\leq\lfloor(N-1)/2\rfloor$ and $f\leq\lfloor(N-1)/3\rfloor$ for synchronous and asynchronous settings respectively, a clear security advantage over strong BFT consensus.
@@
Table 1: Byzantine Agreement (BA) Primitives
Primitive Input Output Asynchronous Byzantine Fault Tolerance
Binary BA  (Atomic Broadcast) Input: One value
Output: Commit/abort
$f\leq\lfloor(N-1)/3\rfloor$
Multi-valued Validated BA Input: Multiple values
Output: All validated
$f\leq\lfloor(N-1)/3\rfloor$
Strong BFT Consensus Input: Multiple values
Output: One
$f\leq\lfloor(N-1)/(m+1)\rfloor$
Validated Strong BFT Consensus Input: Multiple values
Output: One validated
$f\leq\lfloor(N-1)/3\rfloor$
Voting Complexity Comparison in a Nutshell: Nodes in Binary BA, Strong BFT consensus and Validated Strong BFT consensus model makes only one vote per voting round, until the protocol terminated. In contrast, MVBA requires nodes to find all validated input values and vote on each and everyone of them for their validity, resulting in much higher voting complexity.
@@ 3.1 Key Elements of the Protocol
To help understand our proposal, we first highlight four key aspects of our protocol that depart from existing asynchronous protocols for achieving strong BFT consensus:
1). Periodic Leader-based Block Proposal and Allowing Unknown $m$: The classical protocol for strong BFT consensus (see Appx. A.1) doesn’t specify the origin of the $m$ input values. Implementing it in an asynchronous leader-based blockchain would require $m\geq N-f$ nodes to propose blocks. In contrast, our protocol uses a single periodically rotating leader node to propose a block in each time window. However, a block can arrive at any time due to arbitrary but finite message delays. We guarantee the correctness of the consensus protocol even with inconsistencies of blocks among nodes and new blocks being discovered during voting rounds.
2). Increased Adversarial Tolerance: The classical protocol tolerates $f\leq\lfloor(N-1)/(m+1)\rfloor$ adversarial nodes, $m>2$. Our protocol allows $f\leq\lfloor(N-1)/3\rfloor$.
3). Forcing Verdict Unification via Voting Rules: A vote is structurally linked to $N-f$ votes of the last round. A partial voting history can be recursively obtained via a vote graph starting from this vote. This structure enables verification of voting rules compliance. Specifically, any recipient of a vote can unambiguously obtain this partial voting history and determine whether the vote was validly cast for a block. In other words, the voter must justify why it voted for a particular block — namely, that it was the most supported block according to the voting history. This design ensures that nodes gradually converge on a single block. Over time, the Byzantine nodes lose the ability to justify voting for alternative blocks.
4). Concurrent Consensus Derivation for Multiple Instances: In contrast to existing BFT protocols for blockchains, which run only one instance at a time, our protocol nests and concurrently executes multiple instances. Specifically, each blockchain height $k\geq 0$ starts a new instance $k$. The voting process is structured such that a vote cast at height $k$ corresponds to a round 0 vote for instance $k$, a round 1 vote for instance $k-1$, and so on, up to a round $k$ vote for instance 0.
@@ 3.2 Definitions
Vote Chain. The vote chain for node $i$ consists of successive votes, where each vote of node $i$ at height $k\geq 1$ has a directed link to node $i$’s vote at height $k-1$.
@@
Vote ($V^{(k)}_{i}$). This is the vote cast by node $i$ for a block at consensus height $k$. For $k\geq 1$, $V^{(k)}_{i}$ is a tuple consisting of the hash value of a block that node $i$ votes for at height $k$ and the hash values of $N-f$ votes from height $k-1$.
@@
Vote View ($\operatorname{VV}_{i}^{(k)}$). This is an array of integers, where $\operatorname{VV}_{i}^{(k)}[B]$ is defined for a block $B$ of height less than $k$. $\operatorname{VV}_{i}^{(k)}[B]$ contains the sum of the votes in $G_{i}^{(k)}$, which were cast on $B$. Note that $\operatorname{VV}_{i}^{(k)}$ does not count $V^{(k)}_{i}$.
Vote Count ($\operatorname{VC}_{i}^{(k)}$). This is an array of integers, where $\operatorname{VC}_{i}^{(k)}[B]$ sums the votes cast on block $B$ and any blocks up until consensus height $k-1$ stemming from $B$, using the data from $\operatorname{VV}_{i}^{(k)}$.
@@ 3.3 Consensus Chain Propagation
For simplicity of explanation, we assume all $N$ nodes are indexed and have access to a local clock [^2] For example, there can be a pre-set agreement for nodes to start the system together at 00:00 UTC on a day., denoted by $T$, which is not required to align perfectly among the nodes. Node 0 generates the Genesis block at $T=0$, which is consensually accepted by all nodes. At the beginning of each predefined period $\Delta BI$, the node with index $i=\lfloor\frac{T}{\Delta BI}\rfloor\mod N$ proposes a block to extend the consensus chain known to node $i$  [^3] Branches (forkings) may exist, resulting in multiple consensus chains. Because of arbitrary but finite network delays, these consensus chains may or may not be known to node $i$.. As nodes are not assumed to be aligned precisely according to $T$, the adversary could effectively propose a block at any time for any branch. Further discussion on this design see Appx. C.2. Fig. 1 shows an example of consensus chain propagation. All blocks stem from the Genesis block and form a tree structure.
Figure 1: An example of consensus chain propagation: Node 1 generated Block 1, stemming from Block 0. Block 2 did not extend Block 1; instead, it extended Block 0. Block 3 extended Block 1, etc. This figure also illustrates the vote view $\operatorname{VV}_{j}^{(k)}$ and the vote count $\operatorname{VC}_{j}^{(k)}$, for $k=3$. The numbers in the blue and red circles on Block X are $\operatorname{VV}_{j}^{(3)}[X]$ and $\operatorname{VC}_{j}^{(3)}[X]$ respectively. $G_{j}^{(3)}$ contains votes until height 2 to the knowledge of node $j$.
@@ 3.4 Voting
All votes at height 0 are for the Genesis block. An honest node $i$ casts a vote at vote height $k>0$ only upon receipt of $N-f$ votes of height $k-1$  [^4] Note that if some of the $N-f$ votes are labeled as non-compliant votes by the node, this node will wait until it receives $N-f$ compliant votes. Sec. 3.4.1 discusses when a vote will be labeled non-compliant., which can be used to construct $G_{i}^{(k)}$ and derive both $\operatorname{VV}_{i}^{(k)}$ and $\operatorname{VC}_{i}^{(k)}$  [^5] Note that $G_{i}^{(k)}$ may contain double-voting votes, which need to be deducted before deriving $\operatorname{VC}_{i}^{(k)}$. See the filtering rule in Sec. 3.4.1..
@@
      Let $B_{k^{\prime}}$ be the block at consensus height $k^{\prime}$ that $B_{k}$ stems from, where $0\leq k^{\prime}<k$, and $B_{0}$ be the Genesis block. $\forall k^{\prime}\in\{1,...,k-1\}$, $B_{k^{\prime}}$ must have the highest vote count among blocks at consensus height $k^{\prime}$ that stem from $B_{k^{\prime}-1}$ according to $\operatorname{VC}_{i}^{(k)}$.
@@
4:         Generate $V_{i}^{(currentHeight)}$ that links to $N-f$ compliant votes of $currentHeight-1$. Build $G_{i}^{(currentHeight)}$ and then calculate $\operatorname{VC}_{i}^{(currentHeight)}$.
5:         Cast $V_{i}^{(currentHeight)}$ on a block of $currentHeight$ from the most supported branch (per criteria. 2 of the honest model, according to $\operatorname{VC}_{i}^{(currentHeight)}$).
@@
Vote reasoning. When receiving $V_{j}^{(k)}$ from any node $j$, any node can acquire the blocks known to node $j$ and their associated votes via $G_{j}^{(k)}$. Fig. 1 shows an example with six blocks up to height 3 and the information derived from $V_{j}^{(3)}$. The numbers in the blue and red circles on block $X$ are $\operatorname{VV}_{j}^{(3)}[X]$ and $\operatorname{VC}_{j}^{(3)}[X]$ respectively. No information about block 4 and block 6 are indicated in $V_{j}^{(3)}$ as $G_{j}^{(3)}$ only contains votes up until height 2. One cannot infer if node $j$ knows any block of height 3 base on $V_{j}^{(3)}$. But, $V_{j}^{(3)}$ must be cast for block 6 as it is in the branch rooted by block 5, which has the highest vote count among blocks of height 2 stemming from block 1. Block 1 has the highest vote count among blocks of height 1 that stems from the Genesis block.
Remarks: The reason for using $\operatorname{VC}$, which sums up the votes of a block’s branch as the vote count of this block, is that once this block is consensually accepted, later consensually accepted blocks must come from its branch. Thus, the goal is to decide on which branch (rooted by a block) is most supported, not merely the most supported block.